i m as much older than my brother who is 10 years as i am younger than my father who is 70 yrs.how old i am?
Answer by ankor@dixie-net.com(22740)

a - 10 = 70 - a
a + a = 70 + 10
2a = 80
a = 40 yrs old (30 yrs old than his brother, 30 yrs younger than father)